A large man sits on a four-legged chair with his feet off the floor. The combined mass of the man and chair is 95.0 kg. If the chair legs are circular and have a radius of 0.500 cm at the bottom, what pressure does each leg exert on the floor?


1
Expert's answer
2021-10-25T09:56:44-0400

"P = \frac{F}{S}"

"m = 95 kg"

"g = 9.8\frac{m}{s^2}"

"F = mg = 95*9.8= 931N"

"r = 0.5cm= 0.5*10^{-2}m"

"n = 4"

"S = n\pi r^2= 4*\pi*0.25*10^{-4}= \pi*10^{-4}m^2"

"P = \frac{F}{S}= \frac{931}{\pi *10^{-4}}\approx2.96*10^6Pa"


"\text{Answer: }2.96*10^6\ Pa"